# Deep learning for dynamic graphs: models and benchmarks

### v1

@@ VI The benchmark problem
TABLE I: A selected list of datasets used in dynamic graphs representation learning field. The “$\mathcal{C}$” in the type column means C-TDG, “$\mathcal{D}$” corresponds to D-TDG, and “$\mathcal{ST}$” to spatio-temporal graph.
@@
VI-A Spatio-temopral graph benchmark
In the spatio-temporal setting, we consider three graph datasets for traffic forecasting, i.e., Metr-LA [37], Montevideo [66], and PeMSBay [37]. Specifically,
We designed each model as a combination of three main components. The first is the encoder which maps the node input features into a latent hidden space; the second is the DGN which computes the spatio-temporal convolution; and the third is a readout that maps the output of the convolution into the output space. The encoder and the readout are MLPs that share the same architecture among all models in the experiments. We performed hyper-parameter tuning via grid search, optimizing the Mean Absolute Error (MAE). We perform a time-based split of the dataset which reserves the first 70% of the data as training set, 15% of the following data as validation set, and the last 15% as test set. We trained the models using Adam optimizer for a maximum of 1000 epochs and early stopping with patience of 50 epochs on the validation error. For each model configuration, we performed 5 training runs with different weight initialization and report the average of the results. We report in Table II the grid of hyper-parameters exploited for this experiment.
TABLE II: The grid of hyper-parameters employed during model selection for the spatio-temporal tasks.
Configs
learning rate $10^{-2}$, $10^{-3}$, $10^{-4}$
weight decay $10^{-3}$, $10^{-4}$
embedding dim 1, 2, 4, 8
$\sigma$ ReLU
Chebishev poly. filter size 1, 2, 3
Results
In Table III we report the results on the spatio-temporal-based experiments.
TABLE III: Mean test scores of the spatio-temporal models and std averaged over 5 random weight initializations. MAE is the optimized metric. The lower, the better.
Montevideo Metr-LA PeMSBay
Model MAE MSE MAE MSE MAE MSE
A3TGCN $0.3962_{\pm 0.0021}$ $1.0416_{\pm 0.0047}$ $0.3401_{\pm 0.0008}$ $0.3893_{\pm 0.0039}$ $0.2203_{\pm 0.0105}$ $0.2540_{\pm 0.0039}$
DRCNN $0.3499_{\pm 0.0006}$ $1.0686_{\pm 0.0012}$ 0.1218±${}_{\pm 0.0013}$ $0.0960_{\pm 0.0017}$ 0.0569±${}_{\pm 0.0004}$ $0.0398_{\pm 0.0002}$
GCRN-GRU 0.3481±${}_{\pm 0.0008}$ $1.0534_{\pm 0.0062}$ $0.1219_{\pm 0.0007}$ $0.0973_{\pm 0.0009}$ $0.0571_{\pm 0.0005}$ $0.0404_{\pm 0.0007}$
GCRN-LSTM $0.3486_{\pm 0.0026}$ $1.0451_{\pm 0.0099}$ $0.1235_{\pm 0.0009}$ $0.0985_{\pm 0.0004}$ $0.0593_{\pm 0.0004}$ $0.0436_{\pm 0.0001}$
TGCN $0.4024_{\pm 0.0022}$ $1.0678_{\pm 0.0049}$ $0.3422_{\pm 0.0046}$ $0.3891_{\pm 0.0058}$ $0.2109_{\pm 0.0039}$ $0.2466_{\pm 0.0033}$
Overall, DCRNN and GCRN-GRU achieve the better performance on the selected tasks. Interestingly, they both rely on Chebyshev spectral convolution and GRU, but with different architectural structure. Indeed, DCRNN employs a stacked architecture, while GCRN-GRU embeds the DGN into the RNN, enabling a combined modeling of the temporal and spatial information. This result shows that there is not a superior architectural design, in these tasks. However, it seems relevant to include a bigger neighborhood in the computation (e.g., by exploiting a larger Chebishev polynomial filter size). Indeed, even if A3TGCN employs an attention mechanism to capture more global information, it is not enough to achieve comparable performance to DCRNN or GCRN-based approaches.
VI-B D-TDG benchmark
  • Bitcoin-$\alpha$ [70, 71]: a who-trusts-whom network of bitcoin users trading on the platform http://www.bitcoin-alpha.com.
We performed hyper-parameter tuning via grid search, optimizing the Mean Absolute Error (MAE) in the case of node regression and Area Under the ROC curve (AUROC) in the case of node classification and link prediction. We considered the same experimental setting, split strategy, and architectural choice as for the spatio-temporal graphs. In the case of link prediction, we perform negative sampling by randomly sampling non-occurring links from the next future snapshots. We note that in the case of DynGESN, the model employs fixed and randomized weights and only the final readout is trained. We report in Table IV the grid of hyper-parameters exploited for this experiment.
TABLE IV: The grid of hyper-parameters employed during model selection for the D-TDG tasks. The “***” value refer only to LRGCN model, while $\diamond$” to DynGESN.
Configs
learning rate $10^{-2}$, $10^{-3}$, $10^{-4}$
weight decay $10^{-3}$, $10^{-4}$
embedding dim 8, 16, 32
$\sigma$ ReLU
Chebishev poly. filter size 1, 2, 3
n. bases in the basis-decomposition regularization $*{}^{*}$ None, 10, 30
leakage (i.e., $\gamma$${}^{\diamond}$ 0.1, 0.5, 0.9
random weight init. ${}^{\diamond}$ 0.1, 0.5, 0.9
Results
Table V shows the results on general D-TDGs. Differently than the spatio-temporal setting, different tasks benefit from different architectures. Indeed, integrating topology’s changes (such as in GCLSTM and LRGCN) is more effective in link prediction tasks, while evolving the parameters of the DGN is more beneficial for node-level tasks, since it is more difficult to change the parameters of a static DGN to predict the topological evolution of the system. Notably, DynGESN achieves comparable results by exploiting only few trainable parameters, showing an advantageous tradeoff between performance and complexity. This makes it an ideal choice when the computational resources are limited.
@@ Results
TABLE V: Mean test scores and std of DGNs for general D-TDGs averaged over 5 random weight initializations. For MAE and MSE scores, lower values corresponds to better performances, while for AUROC and F1 the higher values are better. MSE and AUROC are the optimized metrics.
Node-level tasks Link-level tasks
@@
Model MAE MSE AUROC F1 AUROC F1 AUROC F1
DynGESN $0.1944_{\pm 0.0056}$ $0.3708_{\pm 0.0411}$ $50.57_{\pm 1.34}$ $77.36_{\pm 20.31}$ $86.88_{\pm 0.26}$ $86.02_{\pm 0.35}$ $90.08_{\pm 3.19}$ $89.14_{\pm 3.86}$
EvolveGCN-H $0.1735_{\pm 0.0007}$ $0.2858_{\pm 0.0074}$ $97.32_{\pm 0.08}$ $98.64_{\pm 0.04}$ $63.87_{\pm 16.99}$ $46.57_{\pm 38.49}$ $52.84_{\pm 3.48}$ $10.15_{\pm 12.43}$
EvolveGCN-O $0.1749_{\pm 0.0007}$ $0.3020_{\pm 0.0111}$ 97.47±${}_{\pm 0.01}$ $98.72_{\pm 0.006}$ $57.51_{\pm 15.02}$ $44.22_{\pm 36.92}$ $51.42_{\pm 2.84}$ $45.07_{\pm 27.64}$
GCLSTM 0.1686±${}_{\pm 0.0015}$ $0.2588_{\pm 0.0049}$ $97.24_{\pm 0.20}$ $98.60_{\pm 0.10}$ 91.49±${}_{\pm 0.09}$ $91.62_{\pm 0.11}$ $91.94_{\pm 0.19}$ $91.65_{\pm 0.19}$
LRGCN $0.1693_{\pm 0.0014}$ $0.2507_{\pm 0.0057}$ $97.15_{\pm 0.58}$ $98.55_{\pm 0.30}$ $88.80_{\pm 1.22}$ $89.23_{\pm 1.25}$ 93.14±${}_{\pm 0.52}$ $92.88_{\pm 0.61}$
VI-C C-TDG benchmark
We report in Table VI the grid of hyper-parameters exploited for this experiment.
TABLE VI: The grid of hyper-parameters employed during model selection for the C-TDG tasks.
Configs
learning rate $10^{-3}$,$10^{-4}$
weight decay $10^{-4}$,$10^{-5}$
n. DGN layers 1, 3
embedding dim 32, 64, 96
$\sigma$ tanh
Neighborhood sampler size 5
Results
We report the results of the C-TDG experiments in Table VII.
TABLE VII: Mean test scores and std of DGNs for C-TDGs averaged over 5 random weight initializations. The higher, the better. The models are trained to maximize the AUROC score.
@@
Model AUROC F1 AUROC F1 AUROC F1
EdgeBank 91.82 91.09 96.42 96.29 94.72 94.43
Overall, TGN generally outperforms all the other methods, showing consistent improvements over DyRep and JODIE. This result shows how the spatial information is fundamental for the effective resolution of the tasks. Indeed, an advantage of TGAT and TGN is that they can exploit bigger neighborhoods with respect to DyRep, which uses the information coming from one-hop distance, and JODIE, which only encode the source and destination information. Despite these results, we observe that the temporal information is still extremely relevant to achieve good performance. In fact, the EdgeBank baseline is able to exceed 91% AUROC score by only looking at the graph’s history. This is even more evident in the LastFM task, which, as observed in [72], contains more reoccurring edges with respect to Wikipedia and Reddit. Consequently, such a task is comparatively easier to solve by solely exploiting these temporal patterns. Considering that EdgeBank’s performance is directly correlated to the number of memorized edges, in this task, it is able to outperform all the other methods.
VII Conclusions
In order to further improve the maturity of representation learning for dynamic graphs, we believe that certain aspects should be deepened. A future interesting direction, in this sense, is to extend the work that has been done for heterophilic (static) graphs [73, 74, 75] to the temporal domain. This will require addressing the problem of generating information-rich node representations when neighboring nodes tend to belong to different classes. A similar challenge is the one of heterogeneus dynamic graphs, which contain different types of nodes and links. In this scenario, new architectures should learn the semantic-level information coming from node and edge types, in addition to topological and label information. Although these are interesting future directions, we observe that there are more challenges to be addressed in the future, such as the study in the temporal domain of over-smoothing [76], which is a phenomenon where all node features become almost indistinguishable after few embedding updates; over-squashing [77, 78], which prevents DGNs to propagate and preserve long-term dependencies between nodes; and expressive power of DGNs [23, 79].
Acknowledgements

### v2

@@ Acknowledgements
TABLE VII: A selected list of datasets used in dynamic graphs representation learning field. The “$\mathcal{C}$” in the type column means C-TDG, “$\mathcal{D}$” corresponds to D-TDG, and “$\mathcal{ST}$” to spatio-temporal graph.
@@
TABLE VIII: An overview of the examined models and the specific changes in the graph structure that each model was designed to address, i.e., node additions/deletions and edge addition/deletions. The “$\mathcal{C}$” in the type column means C-TDG, “$\mathcal{D}$” corresponds to D-TDG, and “$\mathcal{ST}$” to spatio-temporal graph.
Node Edge
Name Cit. Type Addition Deletion Addition Deletion
A3TGCN [41] $\mathcal{ST}$
ASTGCN [44] $\mathcal{ST}$
CAW [66] $\mathcal{C}$
CTDNG [65] $\mathcal{C}$
DCRNN [39] $\mathcal{ST}$
DyGrAE [58] $\mathcal{D}$
DyRep [11] $\mathcal{C}$
DynGEM [61] $\mathcal{D}$
DynGESN [48] $\mathcal{D}$
DynGraph2Vec [60] $\mathcal{D}$
E-GCN [57] $\mathcal{D}$
Evolve2Vec [62] $\mathcal{D}$
GC-LSTM [45] $\mathcal{D}$
GCRN [36] $\mathcal{ST}$
JODIE [63] $\mathcal{C}$
LRGCN [46] $\mathcal{D}$
MPNN-LSTM [50] $\mathcal{D}$
NeurTW [67] $\mathcal{C}$
PINT [68] $\mathcal{C}$
ROLAND [51] $\mathcal{D}$
SGP [53] $\mathcal{D}$
STGCN [42] $\mathcal{ST}$
StreamGNN [64] $\mathcal{C}$
T-GCN [9] $\mathcal{ST}$
TGAT [12] $\mathcal{C}$
TGN [10] $\mathcal{C}$
TABLE IX: Comparative analysis with previous benchmarking studies and surveys, assessing the provision of datasets and benchmarks and delineating the types of analyzed graphs. The “$\mathcal{C}$” in the columns means C-TDG, “$\mathcal{D}$” corresponds to D-TDG, and “$\mathcal{ST}$” to spatio-temporal graph.
Survey Year of the last surveyed method Datasets Dynamic graph benchmark Static $\mathcal{ST}$ $\mathcal{D}$ $\mathcal{C}$ static dynamic $\mathcal{ST}$ $\mathcal{D}$ $\mathcal{C}$ Hamilton et al. [93] 2017 Bacciu et al. [7] 2020 Wu et al. [8] 2019 Kazemi et al. [13] 2020 Jiang and Luo [14] 2022 ($\mathcal{ST}$ only) Our 2023
-B Selected Hyper-parameters
To favour reproducibility of the results and transparency of our benchmarking analysis, in Table X, Table XI, and Table XII we report the hyper-parameters selected for benchmarking task and model assessed in our empirical analysis.
TABLE X: The table of selected hyper-parameters for the spatio-temporal tasks.
embedding dim lr weight decay filter size normalization
Montevideo A3TGCN 8 $10^{-2}$ $10^{-3}$
DCRNN 8 $10^{-2}$ $10^{-3}$ 1 -
GCRN-GRU 8 $10^{-2}$ $10^{-3}$ 3 $\mathbf{D}-\mathbf{A}$
GCRN-LSTM 8 $10^{-2}$ $10^{-4}$ 3 $\mathbf{D}-\mathbf{A}$
TGCN 8 $10^{-2}$ $10^{-3}$
Metr-LA A3TGCN 8 $10^{-3}$ $10^{-3}$
DCRNN 8 $10^{-3}$ $10^{-3}$ 2 -
GCRN-GRU 8 $10^{-3}$ $10^{-3}$ 3 $\mathbf{D}-\mathbf{A}$
GCRN-LSTM 8 $10^{-3}$ $10^{-3}$ 3 $\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$
TGCN 8 $10^{-3}$ $10^{-4}$
PeMSBay A3TGCN 8 $10^{-3}$ $10^{-4}$
DCRNN 8 $10^{-4}$ $10^{-4}$ 2 -
GCRN-GRU 8 $10^{-4}$ $10^{-4}$ 2 $\mathbf{D}-\mathbf{A}$
GCRN-LSTM 8 $10^{-4}$ $10^{-4}$ 3 $\mathbf{D}^{-1}\mathbf{A}$
TGCN 8 $10^{-4}$ $10^{-4}$
Traffic A3TGCN 8 $10^{-2}$ $10^{-4}$
DCRNN 8 $10^{-3}$ $10^{-3}$ 3 -
GCRN-GRU 8 $10^{-2}$ $10^{-4}$ 2 $\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$
GCRN-LSTM 8 $10^{-2}$ $10^{-3}$ 2 $\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$
TGCN 8 $10^{-2}$ $10^{-4}$
TABLE XI: The table of selected hyper-parameters for the D-TDG tasks.
embedding
dim
lr
weight
decay
n. bases K normalization
random weight
init. value
$\gamma$
Twitter tennis DynGESN 32 $10^{-2}$ $10^{-3}$ 0.9 0.5
EvolveGCN-H 8 $10^{-2}$ $10^{-3}$ $\mathbf{D}-\mathbf{A}$
EvolveGCN-O 32 $10^{-3}$ $10^{-3}$ $\mathbf{D}-\mathbf{A}$
GCLSTM 32 $10^{-2}$ $10^{-4}$ 2 $\mathbf{D}-\mathbf{A}$
LRGCN 32 $10^{-2}$ $10^{-4}$ None - - - -
Elliptic DynGESN 8 $10^{-2}$ $10^{-4}$ 0.1 0.9
EvolveGCN-H 32 $10^{-4}$ $10^{-4}$ $\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$
EvolveGCN-O 16 $10^{-4}$ $10^{-3}$ $\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$
GCLSTM 8 $10^{-4}$ $10^{-4}$ 1 $\mathbf{D}-\mathbf{A}$
LRGCN 8 $10^{-4}$ $10^{-4}$ 10 - - - -
AS-733 DynGESN 32 $10^{-2}$ $10^{-4}$ 0.9 0.5
EvolveGCN-H 32 $10^{-3}$ $10^{-3}$ $\mathbf{D}-\mathbf{A}$
EvolveGCN-O 16 $10^{-3}$ $10^{-4}$ $\mathbf{D}-\mathbf{A}$
GCLSTM 32 $10^{-2}$ $10^{-3}$ 2 $\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$
LRGCN 32 $10^{-3}$ $10^{-3}$ 10 - - - -
Bitcoin $\alpha$ DynGESN 32 $10^{-2}$ $10^{-4}$ 0.9 0.1
EvolveGCN-H 32 $10^{-4}$ $10^{-3}$ $\mathbf{D}-\mathbf{A}$
EvolveGCN-O 16 $10^{-3}$ $10^{-4}$ $\mathbf{D}-\mathbf{A}$
GCLSTM 16 $10^{-2}$ $10^{-4}$ 3 $\mathbf{D}-\mathbf{A}$
LRGCN 32 $10^{-2}$ $10^{-4}$ 10 - - - -
TABLE XII: The table of selected hyper-parameters for the C-TDG tasks.
embedding dim lr weight decay n. DGN layers DGN dim
Wikipedia DyRep 96 $10^{-3}$ $10^{-4}$
JODIE 96 $10^{-4}$ $10^{-5}$
TGAT 96 $10^{-3}$ $10^{-4}$ 3 96
TGN 96 $10^{-3}$ $10^{-5}$ 3 48
Reddit DyRep 96 $10^{-4}$ $10^{-4}$
JODIE 96 $10^{-4}$ $10^{-5}$
TGAT 96 $10^{-3}$ $10^{-4}$ 3 96
TGN 96 $10^{-3}$ $10^{-5}$ 1 96
LastFM DyRep 96 $10^{-4}$ $10^{-4}$
JODIE 32 $10^{-4}$ $10^{-5}$
TGAT 48 $10^{-4}$ $10^{-4}$ 3 48
TGN 32 $10^{-3}$ $10^{-5}$ 3 16
-C Complete results
Table XIII and Table XIV provide the full results for the discrete and continuous time dynamic graph tasks, including balanced accuracy as an additional metric (with respect to those in the main body of the work). Table XV, Table XVI, and Table XVII, instead, report the minimum and maximum standard deviation (std) of validation scores obtained by fixing individual hyper-parameter values in spatio-temporal, D-TDGs, and C-TDGs, respectively. Thus, we study the stability of training under various hyper-parameters. We observe that models typically exhibit stability across diverse hyper-parameter settings, and that weight decay is among the hyper-parameters with less influence on the final score, while the embedding dimension and learning rate are the most affecting ones. Lastly, we report the average time per forward pass on the whole dataset (measured in seconds) for each model in every task in Table XVIII, Table XIX, and Table XX.
@@
TABLE XIII: Mean test scores and std of DGNs for general D-TDGs averaged over 5 random weight initializations. For MAE and MSE scores, lower values corresponds to better performances, while for Balacend Acc, AUROC and F1 the higher values are better. In red is reported the optimized metric.
Node-level tasks
Twitter tennis Elliptic
MAE MSE AUROC F1 Balanced Acc
DynGESN 0.1944±${}_{\pm 0.0056}$ 0.3708±${}_{\pm 0.0411}$ 51.12±${}_{\pm 1.30}$ 79.2±${}_{\pm 19.62}$ 50.56±${}_{\pm 1.10}$
EvolveGCN-H 0.1735±${}_{\pm 0.0007}$ 0.2858±${}_{\pm 0.0074}$ 48.43±${}_{\pm 2.71}$ 92.54±${}_{\pm 8.39}$ 49.52±${}_{\pm 1.55}$
EvolveGCN-O 0.1749±${}_{\pm 0.0007}$ 0.3020±${}_{\pm 0.0111}$ 45.11±${}_{\pm 1.68}$ 90.80±${}_{\pm 12.67}$ 49.23±${}_{\pm 1.03}$
GCLSTM 0.1686±${}_{\pm 0.0015}$ 0.2588±${}_{\pm 0.0049}$ 45.77±${}_{\pm 1.60}$ 70.84±${}_{\pm 30.01}$ 48.20±${}_{\pm 1.80}$
LRGCN 0.1693±${}_{\pm 0.0014}$ 0.2507±${}_{\pm 0.0057}$ 45.82±${}_{\pm 3.81}$ 65.69±${}_{\pm 20.21}$ 47.84±${}_{\pm 3.37}$
Link-level tasks
AS-773 Bitcoin $\alpha$
AUROC F1 Balanced Acc AUROC F1 Balanced Acc
DynGESN 95.34±${}_{\pm 0.04}$ 79.83±${}_{\pm 5.27}$ 82.80±${}_{\pm 3.40}$ 97.68±${}_{\pm 0.12}$ 69.98±${}_{\pm 1.57}$ 76.79±${}_{\pm 0.93}$
EvolveGCN-H 59.52±${}_{\pm 17.53}$ 39.85±${}_{\pm 34.24}$ 53.72±${}_{\pm 16.79}$ 51.35±${}_{\pm 2.88}$ 29.55±${}_{\pm 30.58}$ 50.69±${}_{\pm 1.69}$
EvolveGCN-O 58.90±${}_{\pm 17.80}$ 29.99±${}_{\pm 37.10}$ 56.99±${}_{\pm 13.97}$ 51.42±${}_{\pm 2.84}$ 31.74±${}_{\pm 29.98}$ 51.42±${}_{\pm 2.84}$
GCLSTM 96.35±${}_{\pm 0.01}$ 91.22±${}_{\pm 0.13}$ 91.11±${}_{\pm 0.06}$ 97.75±${}_{\pm 0.17}$ 91.22±${}_{\pm 1.38}$ 91.72±${}_{\pm 1.11}$
LRGCN 94.77±${}_{\pm 0.23}$ 89.59±${}_{\pm 0.33}$ 89.07±${}_{\pm 0.34}$ 98.05±${}_{\pm 0.03}$ 91.33±${}_{\pm 0.08}$ 91.89±${}_{\pm 0.07}$
TABLE XIV: Mean test scores and std of DGNs for C-TDGs averaged over 5 random weight initializations. The higher, the better. The models are trained to maximize the AUROC score.
Wikipedia Reddit LastFM
AUROC F1 Acc AUROC F1 Acc AUROC F1 Acc
EdgeBank 91.82 91.09 91.82 96.42 96.29 96.42 94.72 94.43 94.72
DyRep 89.72±${}_{\pm 0.59}$ 79.02±${}_{\pm 0.91}$ 80.46±${}_{\pm 0.63}$ 97.69±${}_{\pm 0.04}$ 92.12±${}_{\pm 0.13}$ 92.02±${}_{\pm 0.19}$ 78.41±${}_{\pm 0.50}$ 71.80±${}_{\pm 0.92}$ 68.63±${}_{\pm 1.01}$
JODIE 94.94±${}_{\pm 0.48}$ 87.52±${}_{\pm 0.39}$ 87.85±${}_{\pm 0.44}$ 96.72±${}_{\pm 0.21}$ 89.97±${}_{\pm 0.66}$ 89.48±${}_{\pm 0.86}$ 69.32±${}_{\pm 1.33}$ 63.95±${}_{\pm 2.64}$ 62.08±${}_{\pm 2.78}$
TGAT 95.54±${}_{\pm 0.22}$ 88.11±${}_{\pm 0.45}$ 88.58±${}_{\pm 0.31}$ 98.41±${}_{\pm 0.01}$ 93.58±${}_{\pm 0.05}$ 93.63±${}_{\pm 0.04}$ 81.97±${}_{\pm 0.08}$ 70.96±${}_{\pm 0.24}$ 72.64±${}_{\pm 0.09}$
TGN 97.07±${}_{\pm 0.15}$ 90.49±${}_{\pm 0.24}$ 90.66±${}_{\pm 0.22}$ 98.66±${}_{\pm 0.04}$ 94.20±${}_{\pm 0.15}$ 94.19±${}_{\pm 0.17}$ 79.84±${}_{\pm 1.58}$ 71.09±${}_{\pm 2.36}$ 63.13±${}_{\pm 7.03}$
TABLE XV: The minimum and maximum standard deviation of validation scores obtained by fixing individual hyperparameter values in spatio-temporal models. The hyperparameter names corresponding to these values are also provided for reference. “wd” means weight decay, “ed” embedding dimension, “lr” learning rate, and “fs” filter size.
Montevideo MetrLA PeMSBay Traffic min max min max min max min max A3TGCN ed: 0.009 lr: 0.015 wd: 0.099 ed: 0.127 ed: 0.039 lr: 0.081 ed: 0.006 lr: 0.009 DCRNN ed: 0.016 lr: 0.025 wd: 0.019 ed: 0.034 ed: 0.050 fs: 0.131 wd: 0.017 ed: 0.020 GCRN-GRU lr: 0.038 K: 0.043 wd: 0.057 K: 0.071 ed: 0.086 K: 0.150 wd: 0.040 K: 0.047 GCRN-LSTM wd: 0.031 ed: 0.039 wd: 0.056 ed: 0.082 ed: 0.072 K: 0.155 wd: 0.024 lr: 0.031 TGCN ed: 0.013 lr: 0.018 wd: 0.099 ed: 0.120 ed: 0.041 lr: 0.087 ed: 0.006 lr: 0.009
TABLE XVI: The minimum and maximum standard deviation of validation scores obtained by fixing individual hyperparameter values in D-TDGs models. The hyperparameter names corresponding to these values are also provided for reference. “wd” means weight decay, “ed” embedding dimension, “lr” learning rate, “ns” normalization scheme, and “nb” number of bases.
Twitter tennis Elliptic AS-733 Bitcoin-$\alpha$ min max min max min max min max DynGESN lr: 0.004 $\sigma$: 0.006 lr: 0.004 ed: 0.005 ed: 0.003 $\sigma$: 0.005 $\gamma$: 0.009 lr: 0.052 EvolveGCN-H wd: 0.011 lr: 0.012 lr: 0.003 ed: 0.005 wd: 0.082 ed: 0.109 ns: 0.018 ed: 0.020 EvolveGCN-O ed: 0.010 lr: 0.012 lr: 0.005 ed: 0.007 ed: 0.095 wd: 0.105 lr: 0.022 wd: 0.024 GCLSTM lr: 0.012 ed: 0.016 wd: 0.004 ed: 0.005 K: 0.201 ed: 0.226 K: 0.025 lr: 0.207 LRGCN nb: 0.008 ed: 0.01 wd: 0.005 nb: 0.005 nb: 0.036 lr: 0.057 nb: 0.003 ed: 0.004
TABLE XVII: The minimum and maximum standard deviation of validation scores obtained by fixing individual hyperparameter values in C-TDGs models. The hyperparameter names corresponding to these values are also provided for reference. “wd” means weight decay, “ed” embedding dimension, “lr” learning rate, “gl” number of DGN layers, “ge” DGN dimension, and “re” readout embedding dim.
Wikipedia Reddit LastFM min max min max min max DyRep ed: 0.004 wd: 0.009 lr: 0.002 wd: 0.003 lr: 0.004 ed: 0.007 JODIE lr: 0.023 ed: 0.027 wd: 0.007 ed: 0.007 wd: 0.004 ed: 0.006 TGAT gl: 0.004 re: 0.023 ge: 0.002 re: 0.011 gl: 0.007 wd: 0.030 TGN ed: 0.004 wd: 0.005 wd: 0.004 ge: 0.006 gl: 0.012 lr: 0.033
TABLE XVIII: Average time to execute a forward pass on the whole dataset (measured in seconds) and std of the best configuration of each model in each task in the spatio-temporal setting, averaged over 5 repetitions. The evaluation were carried out on an Intel(R) Xeon(R) Gold 6240R CPU @ 2.40GHz.
Montevideo Metr-LA PeMSBay Traffic
A3TGCN 1.59±${}_{\pm 0.05}$ 95.12±${}_{\pm 0.53}$ 167.05±${}_{\pm 1.96}$ 18.81±${}_{\pm 1.04}$
DCRNN 3.15±${}_{\pm 0.1}$ 291.44±${}_{\pm 0.76}$ 366.34±${}_{\pm 2.73}$ 112.49±${}_{\pm 11.52}$
GCRN-GRU 4.89±${}_{\pm 0.1}$ 216.99±${}_{\pm 3.77}$ 289.38±${}_{\pm 3.38}$ 32.44±${}_{\pm 1.03}$
GCRN-LSTM 6.92±${}_{\pm 0.13}$ 313.27±${}_{\pm 4.09}$ 534.65±${}_{\pm 5.85}$ 46.08±${}_{\pm 1.6}$
TGCN 1.64±${}_{\pm 0.07}$ 95.08±${}_{\pm 1.43}$ 165.19±${}_{\pm 2.69}$ 17.96±${}_{\pm 0.52}$
TABLE XIX: Average time to execute a forward pass on the whole dataset (measured in seconds) and std of the best configuration of each model in each task in the generic D-TDG, averaged over 5 repetitions. The evaluation were carried out on an Intel(R) Xeon(R) Gold 6240R CPU @ 2.40GHz.
@@
DynGESN 0.04±${}_{\pm 2\cdot 10^{-3}}$ 0.03±${}_{\pm 0.01}$ 0.53±${}_{\pm 0.13}$ 0.15±${}_{\pm 0.01}$
EvolveGCN-H 0.38±${}_{\pm 0.01}$ 21.14±${}_{\pm 0.07}$ 3.73±${}_{\pm 0.18}$ 2.33±${}_{\pm 0.03}$
EvolveGCN-O 0.11±${}_{\pm 0.02}$ 19.5±${}_{\pm 0.25}$ 2.35±${}_{\pm 0.29}$ 1.07±${}_{\pm 0.03}$
GCLSTM 1.08±${}_{\pm 0.42}$ 1.2±${}_{\pm 0.13}$ 31.14±${}_{\pm 0.83}$ 28.37±${}_{\pm 1.59}$
LRGCN 1.63±${}_{\pm 0.08}$ 3.28±${}_{\pm 0.28}$ 21.66±${}_{\pm 2.65}$ 24.77±${}_{\pm 3.21}$
TABLE XX: Average time to execute a forward pass on the whole dataset (measured in seconds) and std of the best configuration of each model in each task in the C-TDG setting, averaged over 5 repetitions. The evaluation were carried out on an Intel(R) Xeon(R) Gold 6278C CPU @ 2.60GHz.
@@
DyRep 13.95±${}_{\pm 1.05}$ 99.11±${}_{\pm 9.04}$ 143.15±${}_{\pm 12.72}$
JODIE 12.66±${}_{\pm 1.98}$ 83.27±${}_{\pm 6.47}$ 117.17±${}_{\pm 6.86}$
TGAT 36.84±${}_{\pm 2.09}$ 303.7±${}_{\pm 7.8}$ 167.65±${}_{\pm 13.33}$
TGN 28.35±${}_{\pm 2.14}$ 114.73±${}_{\pm 14.77}$ 178.15±${}_{\pm 7.87}$